# Leak Proof PDBBind: A Reorganized Dataset of Protein-Ligand Complexes for More Generalizable Binding Affinity Prediction

### v2

@@ 3.4 Rescoring Docked Poses and Pose Selection using the Retrained Scoring Functions
Following a recent study that suggests a modified set of empirical parameters for AutoDock Vina that improves its ranking capabilities, we have also evaluated its performance on the redocked BDB2020+ benchmark set.88 To better separate pose selection capabilities from scoring capabilities, we report metrics on both the redocked structures with the best docking scores, and the redocked poses that have the lowest RMSDs to the cocrystal structures.
Table 4: Performance comparisons on scoring with redocked structures of the BDB2020+ benchmark set in terms of root mean square error (RMSE) and Pearson correlation coefficient (R) for different models before and after retraining using LP-PDBBind. The retrained model with the best performance on the BDB2020+ benchmark dataset was used in this benchmark. AutoDock Vina with the newly suggested weights in Ref. 88 was also added for comparison.
@@
AutoDock Vina with Modified Weights (Ref. 88) 3.35 / 0.25 /
@@
AutoDock Vina with Modified Weights (Ref. 88) 3.40 / 0.17 /
@@
Again, we observed a consistent improvement over the original models after retraining the SFs using LP-PDBBind, which means even though the models were trained using cocrystal structures, they could be used to rescore docked poses and achieve higher accuracies when compared with experimental binding affinities. In the top-scored redocked structure in BDB2020+ benchmark set, AutoDock Vina has improved from 2.12 to 1.99 kcal/mol in RMSE and the correlation coefficients improved from 0.22 to 0.26. Compared to the suggested new weights in Ref.88, the AutoDock Vina scoring function retrained with LP-PDBBind still has better scoring and ranking capabilities. The improvements on the retrained RF-Score and IGN models are also distinct, especially in the correlation coefficients that have improved from 0.34 to 0.49 for RF-Score and from 0.33 to 0.46 for IGN. The performance on the redocked structures that are closest to the known cocrystal structure eliminates the influence of bad docked poses and purely reflects the scoring and ranking capabilities of the retrained models on correctly docked structures. The performance was similar to that evaluated on the best-scored poses, further showing that the improvements obtained by retraining the models using cocrystal structures are transferable to docked structures.
However, we would like to emphasize that the newly trained Vina scoring function is not intended for docking, because a scoring function designed for docking should be trained with high energy decoy poses which are missing in the PDBBind dataset. Consequently, the repulsion parameter in our retrained Vina scoring function (0.014212) is much smaller than the original value (0.840) and the scoring function cannot differentiate bad poses that have atoms too close to each other. Nevertheless, the improvements shown for scoring power and ranking power are meaningful, because the virtual screening process can be broken into multiple steps, and we could use the original Vina SF for docking and the retrained one to rescore top poses to achieve better accuracy.
@@ 4 Discussion and Conclusion
The area of computational drug discovery relies on generalizable scoring functions that have robust scoring and ranking power of binding affinities of ligand-protein complexes. However due to the lack of independently built datasets that test the true generalizability of the SFs, it is hard to differentiate among the plethora of many models which have been trained and tested on the original split of the PDBBind dataset. As we have shown, there is too much overlap between the PDBBind general and refined data used for training with the core subset, leading to the possibility of inflated performance metrics that in turn lead to false confidence in how such models will perform on new protein-ligand complexes.89
In order to reduce data similarity between training, validation, and test data of the PDBBind dataset, we have developed LP-PDBBind using an iterative process to select most similar data first into the test set, and then validation set, so that the final training dataset has low similarity with validation or test dataset. We have also cleaned the PDBBind data in multiple ways: CL1 removed covalent ligand-protein complexes, the low populations of drug molecules with underrepresented chemical elements, and complexes with steric clashes. In addition to CL1, the CL2 level of cleaning aimed for consistent measures of binding free energies by either converting $K_{d}$ or eliminating data reported as $\mathrm{IC_{50}}$. Finally, CL3 eliminated the general set to perform splits on the refined and core set data which is deemed of higher quality. However, it is always a trade-off between the quality of data and the amount of data, and all of the results reported here were based on training on CL1 data and testing on CL2. In addition we have isolated the covalent binding data from PDBBind that may help newly formulated SF predict these type of protein-ligand complexes.
@@
Furthermore, to provide a benchmark dataset truly independent of PDBBind, we have compiled the BDB2020+ dataset derived from the BindingDB database deposited after 2020 to minimize time-cutoff based data leakage, and further ensuring that there is no overlap with PDBBind. Additionally, we also tested ranking power through construction of the Mpro and EFGR ligand-protein complex series. The SARS-CoV-2 Mpro protein has high similarity counterparts in the PDBBind dataset, and the other series involving EGFR does not have anything similar in the training dataset of PDBBind. These new data should also be useful in future evaluation and/or finetuning of any scoring function.
@@
In this work we utilized the new split of PDBBind using CL1 to retrain AutoDock Vina, IGN, RF-Score and DeepDTA and compared the old models with the retrained models on the LP-PDBBind test set, as well as the fully independent BDB2020+ data, Mpro series, and EGFR series. We have demonstrated that using a different splitting of the same dataset leads to significant performance improvements, and the improvements are transferable from using co-crystal structures as input to using docked structures as input. The comparison between the different benchmark results for the SFs can be well explained by the ”difficulty level” of these datasets, and provide insights about the generalizability of various models. We found that well performing models also have more stable ranking results for different ligands towards a protein target. When the target system is similar to data included in the training dataset, the differentiation between models are not obvious. However, when the protein target is not similar to anything in the training dataset, we found that different SFs demonstrate quite different generalization capabilities; the IGN model retrained with LP-PDBBind is recommended due to its reliable good scoring and ranking power.
@@
However, since the PDBBind dataset does not contain decoy structures that are important to allow the models to learn characteristics of true complex structures, the retrained SFs are not expected to have better docking power for recognizing native ligand bound poses. The usefulness of the retrained SF relies on their capability to rescore docked poses and achieve better accuracy when compared to experimental binding affinity measurements. Meanwhile, it would be worthwhile to generate decoy structures for the BDB2020+ dataset to better benchmark the docking and screening powers of the SFs independent of the PDBBind dataset. In summary, the cleaned LP-PDBBind data in its current form is a valuable resource for training more transferable SFs, and we hope reporting evaluation metrics on the BDB2020+ dataset can also become a common practice for future SFs.
@@
We fulsomely acknowledge Charles Brooks for his distinguished career and sustained excellence in biophysics. We sincerely thank Prof. Michael Gilson for assistance with the BDBind data set. This work was supported by National Institute of Allergy and Infectious Disease grant U19-AI171954. This research used computational resources of the National Energy Research Scientific Computing, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.
5 Availability and Implementation
All the code for creating LP-PDBBind dataset, building the BDB2020+ dataset, the prepared dataset files in the .csv format, and python scripts for retraining AutoDock Vina, IGN, RFScore and DeepDTA models are provided in a public accessible GitHub repository: https://github.com/THGLab/LP-PDBBind {suppinfo} Computational procedures and characterization of the PDBBind, EquiBind, and LP-PDBBind data sets, cleaning procedures, and results for original and retrained scoring functions.

### v3

@@ 3.4 Rescoring Docked Poses and Pose Selection using the Retrained Scoring Functions
Following a recent study that suggests a modified set of empirical parameters for AutoDock Vina that improves its ranking capabilities, we have also evaluated its performance on the redocked BDB2020+ benchmark set.55 To better separate pose selection capabilities from scoring capabilities, we report metrics on both the redocked structures with the best docking scores, and the redocked poses that have the lowest RMSDs to the cocrystal structures.
Table 4: Performance comparisons on scoring with redocked structures of the BDB2020+ benchmark. Root mean square error (RMSE) and Pearson correlation coefficient (R) for different models before and after retraining using LP-PDBBind. The retrained model with the best performance on the BDB2020+ benchmark dataset was used in this benchmark. AutoDock Vina with the newly suggested weights in Ref. 55 was also added for comparison.
@@
AutoDock Vina with Modified Weights (Ref. 55) 3.35 / 0.25 /
@@
AutoDock Vina with Modified Weights (Ref. 55) 3.40 / 0.17 /
@@
Again, we observed a consistent improvement over the original models after retraining the SFs using LP-PDBBind, which means even though the models were trained using cocrystal structures, they could be used to rescore docked poses and achieve higher accuracies when compared with experimental binding affinities. In the top-scored redocked structure in BDB2020+ benchmark set, AutoDock Vina has improved from 2.12 to 1.99 kcal/mol in RMSE and the correlation coefficients improved from 0.22 to 0.26. Compared to the suggested new weights in Ref.55, the AutoDock Vina scoring function retrained with LP-PDBBind still has better scoring and ranking capabilities. The improvements on the retrained RF-Score and IGN models are also distinct, especially in the correlation coefficients that have improved from 0.34 to 0.49 for RF-Score and from 0.33 to 0.46 for IGN. The performance on the redocked structures that are closest to the known cocrystal structure eliminates the influence of bad docked poses and purely reflects the scoring and ranking capabilities of the retrained models on correctly docked structures. The performance was similar to that evaluated on the best-scored poses, further showing that the improvements obtained by retraining the models using cocrystal structures are transferable to docked structures.
However, we would like to emphasize that the newly trained Vina scoring function is not intended for docking, because a scoring function designed for docking should be trained with high energy decoy poses which are missing in the PDBBind dataset. Consequently, the repulsion parameter in our retrained Vina scoring function (0.014212) is much smaller than the original value (0.840) and the scoring function cannot differentiate bad poses that have atoms too close to each other. Nevertheless, the improvements shown for scoring power and ranking power are meaningful, because the virtual screening process can be broken into multiple steps, and we could use the original Vina SF for docking and the retrained one to rescore top poses to achieve better accuracy. This is similar to how IGN and RF-Score are used for example, and thus the retrained AutoDock-vina can also be used the same way.
@@ 4 Discussion and Conclusion
The area of computational drug discovery relies on generalizable scoring functions that have robust scoring and ranking power of binding affinities of ligand-protein complexes. However due to the lack of independently built datasets that test the true generalizability of the SFs, it is hard to differentiate among the plethora of many models which have been trained and tested on the original split of the PDBBind dataset. As we have shown, there is too much overlap between the PDBBind general and refined data used for training with the core subset, leading to the possibility of inflated performance metrics that in turn lead to false confidence in how such models will perform on new protein-ligand complexes.6
@@
In order to reduce data similarity between training, validation, and test data of the PDBBind dataset, we have developed LP-PDBBind using an iterative process to select most similar data first into the test set, and then validation set, so that the final training dataset has low similarity with validation or test dataset. Furthermore, to provide a benchmark dataset truly independent of PDBBind, we have compiled the BDB2020+ dataset derived from the BindingDB database deposited after 2020 and ending in 2023 to minimize time-cutoff based data leakage, and further ensuring that there is no overlap with PDBBind. Additionally, we also tested ranking power through construction of the Mpro and EFGR ligand-protein complex series. The SARS-CoV-2 Mpro protein has high similarity counterparts in the PDBBind dataset, and the other series involving EGFR does not have anything similar in the training dataset of PDBBind. These new data should also be useful in future evaluation and/or finetuning of any scoring function.
@@
In this work we utilized the new split of PDBBind to retrain AutoDock Vina, IGN, RF-Score and DeepDTA and compared the old models with the retrained models on the LP-PDBBind test set, as well as the fully independent BDB2020+ data, Mpro series, and EGFR series. We have demonstrated that using a different splitting of the same dataset leads to significant performance improvements, and the improvements are transferable from using co-crystal structures as input to using docked structures as input. The comparison between the different benchmark results for the SFs can be well explained by the ”difficulty level” of these datasets, and provide insights about the generalizability of various models. We found that well performing models also have more stable ranking results for different ligands towards a protein target. When the target system is similar to data included in the training dataset, the differentiation between models are not obvious. However, when the protein target is not similar to anything in the training dataset, we found that different SFs demonstrate quite different generalization capabilities; the IGN model retrained with LP-PDBBind is recommended due to its reliable good scoring and ranking power.
@@
However, since the PDBBind dataset does not contain decoy structures that are important to allow the models to learn characteristics of true complex structures, the retrained SFs are not expected to have better docking power for recognizing native ligand bound poses. The usefulness of the retrained SF relies on their capability to rescore docked poses and achieve better accuracy when compared to experimental binding affinity measurements. Meanwhile, it would be worthwhile to generate decoy structures for the BDB2020+ dataset to better benchmark the docking and screening powers of the SFs independent of the PDBBind dataset. In summary, we have created more transferable SFs, and we hope reporting evaluation metrics on the BDB2020+ dataset can also become a common practice for future SFs.
@@
We sincerely thank Prof. Michael Gilson for assistance with the BDBind data set. This work was supported by National Institute of Allergy and Infectious Disease grant U19-AI171954. This research used computational resources of the National Energy Research Scientific Computing, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.
5 Data Availability and Implementation
All the code for building the BDB2020+ dataset, the prepared dataset files in the .csv format, and python scripts for retraining AutoDock Vina, IGN, RFScore and DeepDTA models are provided in a public accessible GitHub repository: https://github.com/THGLab/LP-PDBBind {suppinfo} Computational procedures and characterization of the PDBBind, EquiBind, and LP-PDBBind data sets, and results for original and retrained scoring functions.
Figure 4: TOC graphic